# &Jiawei Li

### v1

@@ 3.2 Main Results
We present our experimental findings across the above metrics and compare the performance of $\Delta-\mathrm{Influence}$ against baseline methods. Specifically, we report the fraction of detected poisoned samples and the true positive rate in Table 1, and the overall poison success rate along with test accuracy in Figure 3.
Performance of $\Delta$-Influence. As illustrated in Figure 3, with EU (Goel et al., 2023) as the unlearning method, $\Delta$-Influence consistently achieves a poison success rate below 5% across all three types of poisoning attacks and both datasets. This success rate is marked by a , while unsuccessful detections are marked by a $\times$. In contrast, the next best methods, Activation Clustering (ActClust) and EK-FAC, succeed in only 3 out of 6 cases, as highlighted in Table 1. This showcases the substantial improvement in performance of $\Delta-\mathrm{Influence}$.
Among the baseline methods, EK-FAC outperforms ActClust by minimizing the drop in test accuracy, which is indicated by a higher precision in Table 1. Furthermore, $\Delta$-Influence consistently achieves the highest precision, offering better performance with minimal accuracy loss compared to the other methods. Additional experiments detailed in Section 4.1 demonstrate that both label and input augmentations are necessary for $\Delta-\mathrm{Influence}$.
@@
Surprisingly, in the case of the Witches’ Brew attack on CIFAR-10, our $\Delta-\mathrm{Influence}$ method often identifies fewer but a sufficient number of true poisoned samples compared to other methods. We attribute this to the unique behavior of this particular poison. $\Delta-\mathrm{Influence}$ effectively identifies the samples most responsible for the misclassification, and in Witches’ Brew, only a few samples are truly effective for poisoning. Additional experiments in Section 4.2 show that removing the complement of detected poisons does not allow the model to recover, despite the complement set being similar in size or larger.
Conclusion. Overall, $\Delta-\mathrm{Influence}$ offers an effective mechanism for unlearning data poisonining attacks without significantly impacting model performance.
@@
BadNet Frequency Witches’ BadNet Frequency Witches’
@@
Ours (Label-Only) Precision 4.0% 6.3% 1.2% 3.1% 1.1% 0.8%
TPR 97.5% 100% 24.2% 99.1% 100% 73.6%
Ours (Img-Only) Precision 14.4% 31.3% 2.5% 7.6% 0.0% 1.1%
TPR 68.9% 8% 3.2% 50.6% 0.0% 20.8%
Ours (Both) Precision 17.6% 13.3% 3.3% 37.3% 2.9% 2.1%
TPR 99.1% 100% 19.4% 96.9% 100% 62.4%
Table 2: Comparison of Precision and TPR across Label-Only, Image-Only and combined transformation of affected image. Green indicates successful unlearning (poisoning success rate $<5\%$), while red indicates unsuccessful unlearning (See Figure 7 in Appendix 7 for exact poison success rate). Label-only augmentations are highly effective in detecting poisoned samples, whereas image-only augmentations perform poorly. Conversely, image-only augmentations significantly reduce the false positive rate, preserving more clean data and improving detection precision.
4 Unpacking Key Factors in $\Delta$-Infleunce
In this section, we present a series of additional analyses designed to improve the understanding of our proposed $\Delta-\mathrm{Influence}$ method. Specifically, we explore: (i) individual contributions of image and label perturbations, (ii) a counterfactual analysis to determine whether the detected samples are solely responsible for enabling poisoning in the Witches’ Brew attack.
4.1 Perturbing Only Images or Labels
@@
Results. As depicted in Table 2, our ablation study reveals that label-only augmentations achieve high TPR across all poisoning types and datasets, effectively identifying almost all poisoned samples. However, this leads to low precision, resulting in the unnecessary removal of a significant number of clean samples. On the other hand, image-only augmentations exhibit poor TPR, failing at the core task but also rejects lesser number of clean samples (higher precision). In contrast, the combined approach ($\Delta-\mathrm{Influence}$ (Ours)) leverages both label and image perturbations to achieve a balanced performance and successfully detects a high number of poisoned samples (high TPR) while rejecting lesser clean samples (high precision).
Conclusion. Our ablation study underscores the necessity of incorporating both label and image augmentations in the $\Delta-\mathrm{Influence}$ method. Label perturbations are pivotal for enhancing detection accuracy, while image augmentations play a critical role in minimizing false positives.
@@ 4.2 Counterfactual Analysis: Do Detected Samples Account for Poisoning in Witches’ Brew?
Original 62.4% 0% 71.92%
Complement Set 37.6% 100% 72.79%
Table 3: Does the Detected Set Truly Influence the Poison? For Witches’ Brew, we test the “Original" set, representing the poisoned samples identified by $\Delta-\mathrm{Influence}$, and the “Complement" set, which includes all other poisoned samples not detected. The absence of a drop in poison success rate when removing the complement set suggests that the detected set fully captures the poisoning effect. Conversely, removing the detected set completely eliminates the poisoning effect.
Setup. We perform a counterfactual analysis by comparing the original detected set of poisoned samples in the Witches’ Brew attack to its complement set (i.e., all poisoned samples except those detected by $\Delta-\mathrm{Influence}$). This comparison aims to assess whether the detected set exclusively accounts for the poisoning effect.
Results. As presented in Table 3, the removal of the “Original” detected set (19.4% TPR for CIFAR10 and 62.4% TPR for CIFAR100) results in 0% poison success rate, effectively unlearning the poisoning. In stark contrast, removing the “Complement” set (80.6% TPR for CIFAR10 and 37.6% TPR for CIFAR100) maintains a poison success rate of 100%, indicating that the undetected samples do not sufficiently contribute to the poisoning. The complement set achieves higher test accuracy simply because it only contains unaffected samples without false positives.
Conclusion. These results demonstrate that our detected subset accounts for nearly all the poisoning effects in Witches’ Brew, highlighting the unusual nature of the poison and the precision of our $\Delta-\mathrm{Influence}$ algorithm.
5 Investigating Unlearning Methods
This section examines the impact of unlearning methods by evaluating the effectiveness of $\Delta-\mathrm{Influence}$ across various unlearning techniques. Additionally, we investigate how employing a different unlearning algorithm could influences both the success rate of the poisoning attacks and the performance outcomes when applied to various detection methods.
5.1 Which Unlearning Methods Work?
Setup. To evaluate the effectiveness of various unlearning algorithms when paired with our $\Delta-\mathrm{Influence}$ method, we fix the influence function to $\Delta-\mathrm{Influence}$ and vary the unlearning functions. We benchmark several corrective unlearning methods, including EU (Goel et al., 2023), CF  (Goel et al., 2023), SSD  (Foster et al., 2024)SCRUB  (Golatkar et al., 2020), and BadT  (Kurmanji et al., 2023). All methods are implemented using the codebase and training protocols from Goel et al. (2024). Further implementation details, including parameter settings and computational resources, are provided in Appendix B.
Results. As illustrated in Figure 4, our evaluation reveals that CF performs comparably to EU, achieving similar poison removal success rates while offering significant computational gains by avoiding full retraining. CF remains robust against false positives, maintaining high test accuracy. EU effectively removes poisoned samples with no significant drop in test accuracy, albeit at a higher computational cost due to retraining. In contrast, while gradient-ascent-based methods like SCRUB and weight deletion approaches like SSD successfully unlearn poisons, they do so at the expense of model utility due to their susceptibility to false positives. Finally, BadT fails to unlearn poisons effectively.
Conclusion. We recommend EU or CF (Goel et al., 2023) as competitive baselines for corrective unlearning setting using influence functions, and also highlight the importance of robustness towards false positives.
5.2 Does $\Delta$-Influence Perform the Best Across Unlearning Algorithms?
Results. The evaluation results in Figure 5 shows that $\Delta-\mathrm{Influence}$ outperforms other methods, unlearning poisons in all six cases with minimal performance loss. In contrast, EK-FAC, ActClust, and SpecSig performed randomly, achieving unlearning primarily because even a randomly initialized model would not retain poisoning. Performance drops were primarily due to SCRUB’s sensitivity to false positives from its gradient ascent step. FreqDef avoided randomness but failed to unlearn poisons in all cases. Notably, $\Delta-\mathrm{Influence}$ minimized false positives, maintaining consistent and reliable outcomes.
@@ 5.2 Does $\Delta$-Influence Perform the Best Across Unlearning Algorithms?
Figure 6: Scaling to Imagenette. In the top row, results on Imagenette are consistent with previous findings: $\Delta-\mathrm{Influence}$ effectively unlearns all three types of poisons while preserving high test accuracy. In contrast, other detection methods often fail to unlearn or do so at the expense of test accuracy. In the bottom row, EU and CF consistently perform well, with CF (Goel et al., 2023) achieving the smallest accuracy drop and only a minor increase in poison success rate.
@@ 6 Scaling Findings to ImageNette
Setup. To evaluate the scalability and consistency of our $\Delta-\mathrm{Influence}$ algorithm on a more complex and larger dataset, we conduct experiments on Imagenette. The setup is consistent with the experiments in Section 3 with specific adjustments to accommodate Imagenette’s larger image sizes and increased complexity. Specifically, we increase the patch size for BadNet poisoning, use a larger trigger pattern for frequency-based poisoning, and poison a greater fraction of training images (10%). Additionally, for the Witches’ Brew method, we relax the perturbation constraint, setting $\epsilon=32$ instead of $\epsilon=16$.
Results. Replicating our prior experiments on Imagenette, Figure 6 illustrates that $\Delta-\mathrm{Influence}$ continues to achieve the most effective poison unlearning across all attack types, maintaining minimal accuracy loss. Notably, the EK-FAC baseline also successfully unlearns all poisons but incurs a higher false positive rate, leading to significant drops in test accuracy due to the unnecessary removal of clean samples. Additionally, when applying various unlearning algorithms to the samples identified by $\Delta-\mathrm{Influence}$, both CF and EU perform consistently well with CF achieving notably higher accuracy during poison unlearning compared to EU.
@@ 6.1 Scaling to Larger Set of Identified Poisoned Test Samples
Results. We showcase performance in Table 4 for BadNet poison and Table 5 for frequency trigger poison respectively. We observe a consistent trend: as the set of identified poisons increases, the precision improves significantly, leading to a substantial reduction in false positives and ultimately higher test accuracy. Overall, identifying multiple poisoned test points enables more precise detection of poisons in the training set when using $\Delta-\mathrm{Influence}$-like aggregation algorithms across test poisoned points.
@@ 7 Conclusion
In this study, we address a critical issue in corrective machine unlearning: identifying key training samples whose removal can unlearn a data poisoning attack. We address a practical scenario where only a limited number of affected test points are known—potentially discovered post-deployment or through internal testing. To this end, we introduce $\Delta-\mathrm{Influence}$, a novel approach that uses influence functions to trace abnormal model behavior back to the responsible poisoned training data, requiring as little as one affected test example. By retraining without these identified points, $\Delta$-Influence successfully unlearns multiple poisoning attacks across diverse datasets. We evaluate our method against four state-of-the-art detection algorithms and apply five well known unlearning algorithms to the identified training set. Our results demonstrate that $\Delta$-Influence consistently outperforms existing approaches in all tested scenarios. Our findings highlight the potential of influence functions as a foundation for unlearning data poisoning attacks. Additionally, our ablation study sheds light on the strengths and limitations of various poisoning attacks and unlearning algorithms, offering insights that could inform the development of more effective unlearning techniques and robust poisoning attacks for rigorous testing.
@@ Appendix A Connections to Existing Work
Data Poisoning Defences Defenses against data poisoning often involve trigger-pattern reverse engineering using clean data (Wang et al., 2019; Guo et al., 2019; Tao et al., 2022; Dong et al., 2021; Wang et al., 2020). These methods require additional steps such as input pre-filtering, neuron pruning, or fine-tuning (liu_fine-pruning_2018; Chen et al., 2019; Li et al., 2021b; Zeng et al., 2022). Other approaches, like Anti-Backdoor Learning (Li et al., 2021a) and BaDLoss (Alex et al., 2024), necessitate tracking model updates and clean training samples, adding complexity to the defense process.

### v2

@@ 3.2 Main Results
We present our experimental findings across the above metrics and compare the performance of $\Delta-\mathrm{Influence}$ against several baselines. Specifically, we report the precision, TPR, and AuROC of detecting poisons in Table˜1, and the overall PSR and test accuracy after retraining without the identified set in Figure˜3.
Performance of $\Delta$-Influence. As illustrated in Figure˜3, with EU (Goel et al., 2023) as the unlearning method, $\Delta-\mathrm{Influence}$ consistently achieves a poison success rate below 5% across all three types of poisoning attacks and both datasets. This success rate is marked by a , while unsuccessful detections are marked by a $\times$. In contrast, the next best methods, Activation Clustering (ActClust) and EK-FAC, succeed in only 3 out of 6 cases, as highlighted in Table˜1. This showcases the substantial improvement in performance gained by $\Delta-\mathrm{Influence}$. We further validate the effectiveness and robustness of $\Delta-\mathrm{Influence}$ across different unlearning methods, e.g., a popular alternative unlearning algorithm called SCRUB  (Kurmanji et al., 2023) which involves gradient ascent. See detailed results in Figure˜7 in Appendix.
Among the baseline methods, EK-FAC outperforms ActClust by minimizing the drop in test accuracy, also indicated by a higher precision in Table˜1. Furthermore, $\Delta$-Influence consistently achieves the highest precision, offering better performance with minimal accuracy loss compared to the other methods. Additional experiments detailed in Section˜4.1 demonstrate that both label and input augmentations are necessary for $\Delta-\mathrm{Influence}$.
@@
Surprisingly, in the case of the Witches’ Brew attack on CIFAR-10, our $\Delta-\mathrm{Influence}$ method often identifies fewer but a sufficient number of true poisoned samples compared to other methods. We attribute this to the unique behavior of this particular poison. $\Delta-\mathrm{Influence}$ effectively identifies the samples most responsible for the misclassification, and in Witches’ Brew, only a few samples are truly effective for poisoning. Additional experiments in Section˜4.3 show that removing the complement of detected poisons does not allow the model to recover, despite the complement set being similar in size or larger.
Conclusion. Overall, $\Delta-\mathrm{Influence}$ offers an effective mechanism for unlearning data poisonining attacks without significantly impacting model performance. Crucially, it requires no prior knowledge of the attack method, making it more generalizable across various poisoning strategies.
4 Unpacking Key Factors in $\Delta$-Infleunce
We present a series of additional analyses designed to improve the understanding of $\Delta-\mathrm{Influence}$. Specifically, we explore: (i) individual contributions of image and label perturbations, (ii) effectiveness of various unlearning algorithms, (iii) a counterfactual analysis to determine whether the detected samples are solely responsible for enabling poisoning in the Witches’ Brew attack, and (iv) the unreliability of using a known training poison as an attribution target.
4.1 Perturbing Only Images or Labels
@@
Frequency Witches’ BadNet Frequency Witches’ BadNet
@@
Ours (Label-Only) Precision 6.3% 1.2% 4.0% 1.1% 0.8% 3.1%
TPR 100% 24.2% 97.5% 100% 73.6% 99.1%
AuROC 0.94 0.48 0.90 0.92 0.69 0.79
Ours (Img-Only) Precision 28.9% 2.7% 14.4% 0.6% 0.3% 7.6%
TPR 26.4% 13.2% 68.9% 62.4% 40.8% 50.6%
AuROC 0.51 0.25 0.85 0.78 0.62 0.76
Precision 13.3% 3.3% 17.6% 2.9% 2.1% 37.3%
Ours (Both) TPR 100% 19.4% 99.1% 100% 62.4% 96.9%
AuROC 0.96 0.38 0.95 0.96 0.75 0.82
Table 2: Comparison of Precision & TPR & AuROC across Label-Only, Image-Only and combined transformation of affected image. Green indicates successful unlearning (PSR $5\%$), while red indicates unsuccessful unlearning (See Appendix for exact PSR). Label-only augmentations are highly effective in detecting poisoned samples, whereas image-only augmentations perform poorly. Conversely, image-only augmentations significantly reduce the FPR, preserving more clean data and improving detection precision.
@@
Results. As depicted in Table˜2, our ablation study reveals that label-only augmentations achieve high TPR across all poisoning types and datasets, effectively identifying almost all poisoned samples. However, this leads to low precision, resulting in the unnecessary removal of a significant number of clean samples. On the other hand, image-only augmentations exhibit poor TPR, failing at the core task but also rejects less clean samples (higher precision). In contrast, $\Delta-\mathrm{Influence}$ leverages both label and image perturbations to achieve a balanced performance and detects sufficient key poisoned samples while rejecting lesser clean samples (see Figure˜6 in Appendix for detailed unlearning performance).
Conclusion. Our ablation study underscores the necessity of incorporating both label and image augmentations in the $\Delta-\mathrm{Influence}$ method. Label flippings are pivotal for enhancing detection accuracy, while image transformations play a critical role in minimizing false positives.
4.2 Which Unlearning Methods Work?
Setup. To evaluate the effectiveness of various unlearning algorithms when paired with $\Delta-\mathrm{Influence}$, we fix the influence function to $\Delta-\mathrm{Influence}$ and vary the unlearning functions. We benchmark several corrective unlearning methods, including EU (Goel et al., 2023), CF  (Goel et al., 2023), SSD  (Foster et al., 2024)SCRUB  (Kurmanji et al., 2023), and BadT  (Chundawat et al., 2023). All methods are implemented using the codebase and training protocols from Goel et al. (2024). Further implementation details are provided in Appendix B.
Results. As illustrated in Figure˜4, our evaluation reveals that CF performs comparably to EU, achieving similar poison removal success rates while offering significant computational gains by avoiding full retraining. CF remains robust against false positives, maintaining high test accuracy. EU effectively removes poisoned samples with no significant drop in test accuracy, albeit at a higher computational cost due to retraining. In contrast, while gradient-ascent-based methods like SCRUB and weight deletion approaches like SSD successfully unlearn poisons, they do so at the expense of model utility due to their susceptibility to false positives. Finally, BadT fails to unlearn poisons effectively. We recommend EU or CF as competitive baselines for corrective unlearning using influence functions, and also highlight the importance of robustness towards false positives.
Conclusion. We recommend EU or CF as competitive baselines for corrective unlearning using influence functions, and also highlight the importance of robustness towards false positives.
@@ 4.3 Counterfactual Analysis: Do Detected Samples Account for Poisoning in Witches’ Brew?
Setup. The analysis compares the original detected set of poisoned samples in Witches’ Brew to its complement set (i.e., all poisoned samples except those detected by $\Delta-\mathrm{Influence}$). This aims to assess whether the detected set exclusively accounts for the poisoning effect.
Results. As presented in Table˜3, the removal of the “Original” detected set (19.4% TPR for CIFAR10 and 62.4% TPR for CIFAR100) results in 0% poison success rate, effectively unlearning the poisoning. In stark contrast, removing the “Complement” set (80.6% TPR for CIFAR10 and 37.6% TPR for CIFAR100) maintains a poison success rate of 100%, indicating that the undetected samples do not sufficiently contribute to the poisoning. The complement set achieves higher test accuracy simply because it only contains unaffected samples without false positives. These results demonstrate that our detected subset accounts for nearly all the poisoning effects in Witches’ Brew, highlighting the unusual nature of this particular poison as well as the precision of our $\Delta-\mathrm{Influence}$ algorithm.
Conclusion. These results demonstrate that our detected subset accounts for nearly all the poisoning effects in Witches’ Brew, highlighting the unusual nature of this particular poison as well as the precision of our $\Delta-\mathrm{Influence}$ algorithm.
Table 3: Does the Detected Set Truly Influence the Poison? For Witches’ Brew, we test the “Original” set, representing the poisoned samples identified by $\Delta-\mathrm{Influence}$, and the“Complement” set, which includes all other poisoned samples not detected. The absence of a drop in poison success rate when removing the complement set suggests that the detected set fully captures the poisoning effect. Conversely, removing the detected set completely eliminates the poisoning effect.
@@
Original 62.4% 0% 71.9%
Complement Set 37.6% 100% 72.8%
@@
Figure 5: Scaling to Imagenette. In the top row, results on Imagenette are consistent with previous findings: $\Delta-\mathrm{Influence}$ effectively unlearns all three types of poisons while preserving high test accuracy. In contrast, other detection methods often fail to unlearn or do so at the expense of test accuracy. In the bottom row, EU and CF consistently perform well.
4.4 The Causal Pitfall of Targeting a Poisoned Training Sample
Setup. We investigate whether directly using a known poisoned training sample as the attribution target for influence functions can still effectively detect and unlearn poisons (although such availability can be hard to achieve for attacks like Witches’ Brew).
Results. As shown in Table˜4, for CIFAR-10, when the attribution target is a training poison, $\Delta-\mathrm{Influence}$ fails to achieve successful unlearning. This is indicated by the Poison Success Rate remaining at 100%, which signifies that the attack remains fully effective. This inconsistent performance demonstrates that poisoned training samples are unreliable attribution targets for influence-based detection.
Table 4: Failed Unlearning When Targeting a Known Poisoned Training Point. Comparison of using an affected Test point versus a known Train poison as the attribution target.
Identified Point TPR($\uparrow$) Poison Success Rate ($\downarrow$) Test Accuracy ($\uparrow$)
CIFAR10
Test 19.4% 0% 91.0%
Train 8.4% 100% 90.3%
CIFAR100
Test 62.4% 0% 71.9%
Train 84.0% 0% 71.9%
Conclusion. These results justify our choice of using the attribution target from the deployment phase (i.e., test point) instead of training phase. The latter approach yields inconsistent unlearning performance because the causal dependency among poison peers is weaker than the collective influence of all poisons on an affected test point. Our choice is therefore grounded in a causal perspective: the goal is to find training examples responsible for a specific erroneous prediction, making the prediction itself the logical starting point. See Appendix B.1 for a more detailed discussion.
@@ 5 Scaling Findings to ImageNette
Setup. To evaluate the scalability and consistency of our $\Delta-\mathrm{Influence}$ algorithm on a more complex and larger dataset, we conduct experiments on Imagenette. The setup is consistent with the experiments in Section˜3 with specific adjustments to accommodate Imagenette’s larger image sizes and increased complexity. Specifically, we increase the patch size for BadNet poisoning, use a larger trigger pattern for frequency-based poisoning, and poison a greater fraction of training images (10%). Additionally, for the Witches’ Brew method, we relax the perturbation constraint, setting $\epsilon=32$ instead of $\epsilon=16$.
Results. Replicating our prior experiments on Imagenette, Figure˜5 illustrates that $\Delta-\mathrm{Influence}$ continues to achieve the most effective poison unlearning across all attack types, maintaining minimal accuracy loss. Notably, the EK-FAC baseline also successfully unlearns all poisons but incurs a higher false positive rate, leading to significant drops in test accuracy due to the unnecessary removal of clean samples. Additionally, when applying various unlearning algorithms to the samples identified by $\Delta-\mathrm{Influence}$, both CF and EU perform consistently well with CF achieving notably higher accuracy during poison unlearning compared to EU.
@@ 5.1 Scaling to Larger Set of Identified Poisoned Test Samples
Results. We showcase performance in Table˜5 for BadNet poison and Table˜6 for frequency trigger poison respectively. We observe a consistent trend: as the set of identified poisons increases, the precision improves significantly, leading to a substantial reduction in false positives and ultimately higher test accuracy. Overall, identifying multiple poisoned test points enables more precise detection of poisons in the training set when using $\Delta-\mathrm{Influence}$-like aggregation algorithms across test poisoned points.
@@ 6 Conclusion
In this study, we address a critical issue in corrective machine unlearning: identifying key training samples whose removal can unlearn a data poisoning attack. We address a practical scenario where only a limited number of affected test points are known—potentially discovered post-deployment or through internal testing. To this end, we introduce $\Delta-\mathrm{Influence}$, a novel approach that uses influence functions to trace abnormal model behavior back to the responsible poisoned training data, requiring as little as one affected test example. By retraining without these identified points, $\Delta$-Influence successfully unlearns multiple poisoning attacks across diverse datasets. We evaluate our method against five state-of-the-art detection algorithms and apply five well known unlearning algorithms to the identified training set. Our results demonstrate that $\Delta$-Influence consistently outperforms existing approaches in all tested scenarios. Our findings highlight the potential of influence functions as a foundation for unlearning data poisoning attacks. Additionally, our ablation study sheds light on the strengths and limitations of various poisoning attacks and unlearning algorithms, offering insights that could inform the development of more effective unlearning techniques and robust poisoning attacks for rigorous testing.
@@ Appendix A Connections to Existing Work
Data Poisoning Defences Defenses against data poisoning often involve trigger-pattern reverse engineering using clean data (Wang et al., 2019; Guo et al., 2019; Tao et al., 2022; Dong et al., 2021; Wang et al., 2020). These methods require additional steps such as input pre-filtering, neuron pruning, or fine-tuning (Liu et al., 2018; Chen et al., 2019; Li et al., 2021b; Zeng et al., 2022). Other approaches, like Anti-Backdoor Learning (Li et al., 2021a) and BaDLoss (Alex et al., 2024), necessitate tracking model updates and clean training samples, adding complexity to the defense process.